# The Fragile Nature of Road Transportation Systems

### 01 Feb 2024 v1

@@ 5 Numerical simulation
Although researchers always consider MFDs to be well-defined, in the real world, road transportation systems and their MFDs are always subject to uncertainties all the time, as shown in Geroliminis and Daganzo (2008) and in Ambühl et al. (2021) with loop detector data over a year. This uncertainty cannot be reflected in the mathematical analysis, therefore, it is indispensable to show the influence of realistic stochasticity on the performance of transportation systems. Whether the system still maintains the fragile property under real-world uncertainties in MFD.
In this section, we simulate a congestion dissipation process. The MFD of the studied region is created by applying MoC in Daganzo and Geroliminis (2008) with realistic parameters in the city center of Zurich. Some parameters, e.g., free-flow speed, back-propagation speed, maximal density, and capacity are provided in Ambühl et al. (2020) for Zurich with queried routes in Google API and with other validation methods. The total and average lane length for District 1 is queried through SUMO with OpenStreetMap API. The average trip length in the city of Zurich is studied in Schüssler and Axhausen (2008). We introduce stochasticity in District 1 of Zurich with real traffic light data, which is readily available in Genser et al. (2023). Since the signalization in Zurich is actuated based on the present traffic flow, they do not strictly follow a fixed-time signal cycle. Despite this actuation, a concentrated distribution can be easily observed in the dataset and we assume the green split of the cycle follows a normal distribution. The offset is considered to be zero in a way similar to the actuated signal in Yokohama in Daganzo and Geroliminis (2008). The parameters are summarized in Tab. 2.
@@
As the average green time of the signal is 14.8 s and its standard deviation is 2.5 s, following the procedure described in Daganzo and Geroliminis (2008), we can produce three groups of cuts with green time being $\mu_{G}-\sigma_{G}$, $\mu_{G}$, or $\mu_{G}+\sigma_{G}$, and each group of cuts contains at least three cuts, as Fig. 9 shows. The group of cuts with a longer green time of signalization yields a greater MFD, and vice versa. The MFDs share the same maximal vehicle accumulation at around 10000 vehicles.
@@
Now we start the numerical simulation with different initial disruption demands $n_{0}$ from 1000 to 8000 vehicles with a step size of 500 vehicles. The simulation time is the same 5000 seconds for each scenario with a different initial demand. Fig. 10(a) demonstrates that TTS grows exponentially with linearly increasing initial disruption demand. The solid, dashed, or dotted line each represents a deterministic MFD calculated with a signal green time being $\mu_{G}$, $\mu_{G}-\sigma_{G}$, or $\mu_{G}+\sigma_{G}$. Other than these three lines, there are also 500 scattering points forming the blue curve. Each scatter point is composed of a full congestion dissipation process sampled with equal intervals between an initial traffic demand varying from 1000 vehicles to 8000 vehicles. At each time step, a stochastic signal green time is chosen following the normal distribution $G\sim N(14.8,2.5)$ based on real-world data, leading to an uncertain MFD profile as Fig. 10(b) shows. This is an example of how the congestion is dissipated in this urban road network, the green scattering points are sampled from every 80th point from 5000 timesteps in total.
(a) Demand disruption in the real world
@@
Figure 10: Numerical simulation with stochasticity
Since the scattering points closely align with the solid curve, it can be inferred that the influence of realistic stochasticity on the MFD is mostly negligible. However, an intriguing observation is that, when the disruption demand is relatively low, the blue curve dips slightly below the MFD of the solid curve. However, it appears to exceed the TTS of the MFD when the demand is substantial. When we show the distribution of these two curves, as in Fig. 10(c), the curve made from scatter points has a more concentrated distribution at lower TTS while having a greater distribution at higher TTS, indicating a more left-skewed distribution compared to the distribution of TTS without realistic stochasticity. This can also be demonstrated by calculating the skewness of these two curves. When there is no stochasticity, the skewness is 0.50 while the skewness for the curve with stochasticity has a value of 0.53. A greater skewness indicates a more fragile system, which means that by introducing realistic stochasticity, the urban road network becomes even more fragile.
@@ 6 Conclusion
This research systematically demonstrated the fragile nature of road transportation systems with rigorous mathematical analysis and numerical simulation under realistic stochasticity. The mathematical proof comprehends the study of fragility under 1) microscopic - macroscopic, 2) demand disruption - supply disruption, and 3) onset of disruption - recovery from disruption. With essential assumptions regarding the disruption density in comparison to the critical density as well as the base traffic demand, we’ve validated the fragility of road transportation systems from various perspectives. Furthermore, through a numerical simulation with realistic data, we concluded that real-world stochasticity has a limited impact on the fragile characteristics of the system but contributes to rendering the system even more fragile. Lately, researchers have explored the potential utilization of MFD in railway operations (Corman et al., 2019) and in aviation (Safadi et al., 2023). The fragility observed in urban road networks may be extended to various transportation systems. This study aims to offer insights to researchers, emphasizing the fragile characteristics of transportation systems and encouraging the design of antifragile traffic control strategies in the future.

### 04 Mar 2024 v2

@@ 5 Numerical simulation
Even though researchers generally consider MFDs to be well-defined, road transportation systems in the real world and their MFDs are always subject to stochasticity all the time, as shown in Geroliminis and Daganzo (2008); Saffari et al. (2022); Ambühl et al. (2021), and this is the case for FDs as well (Qu et al., 2017; Siqueira et al., 2016). Therefore, when validating a newly proposed traffic control algorithm, it has become a common practice to account for model uncertainties and showcase the method’s robustness, such as in Geroliminis et al. (2013); Haddad and Mirkin (2017); Zhou and Gayah (2023). In our study, however, the model stochasticity cannot be directly reflected in the mathematical analysis, hence, it is indispensable to show the influence of realistic stochasticity on the fragile nature of transportation systems with a numerical simulation, i.e., whether the system still maintains the same fragile response under real-world errors in MFD when a demand or supply disruption is present.
In this section, we simulate the disruption recovery process. The MFD of the studied region is generated by applying MoC following Daganzo and Geroliminis (2008) with realistic parameters in the city center of Zurich. Some parameters, e.g., free-flow speed, back-propagation speed, maximal density, and capacity are provided in Ambühl et al. (2020) for Zurich with queried routes in Google API and with other validation methods. The total and average lane length for the city center is determined through SUMO with OpenStreetMap API. The average trip length of Zurich is studied in Schüssler and Axhausen (2008). We introduce stochasticity in the city center of Zurich with real traffic light data, which is publicized by the Statistical Office of Zurich and accessible in Genser et al. (2023). The authors acknowledge that MoC is developed with the premise of a homogeneous region, and given that the available data is limited to only one main intersection in this region, we assume that this intersection serves as a representative sample for the city center region. Also, since the signalization in Zurich is actuated based on the present traffic flow, they do not strictly follow a fixed-time signal cycle. Despite this actuation, a concentrated distribution can be easily observed in the dataset and we assume the green split of the cycle follows a normal distribution. The offset is considered to be zero in a way similar to the actuated signal in Yokohama in Daganzo and Geroliminis (2008). According to the daily average traffic density of Zurich in Ambühl et al. (2021), we approximate the traffic demand, which is also the trip completion when the traffic state is at equilibrium, is about $0.6veh/s$ for our studied region. This corresponds to an accumulation of around 975 vehicles in the city center. The parameters are summarized in Tab. 2.
@@
Traffic demand
$q$
$\rm veh/s$
0.6
As the average green time of the signal is 14.8 s and its standard deviation is 2.5 s, following the MoC described in Daganzo and Geroliminis (2008), we can produce three groups of cuts with the interval of one standard deviation, i.e., the green time being $\mu_{G}-\sigma_{G}$, $\mu_{G}$, or $\mu_{G}+\sigma_{G}$, and each group of cuts contains at least three cuts, as Fig. 9 shows. The group of cuts with a longer green time of signalization yields a greater MFD, and vice versa. The MFDs share the same maximal vehicle density of around 0.145 veh/m, which corresponds to a maximal accumulation of about 10000 vehicles for the studied region.
@@
Now we start the numerical simulation with different initial disruption demands $n^{\prime}$ from 1000 to 8000 vehicles. The simulation time is 7200 seconds for each scenario with different initial demands. Fig. 10(a) demonstrates that TTS grows exponentially with linearly increasing initial disruption demand, which validates the fragile nature proved with mathematical analysis. The solid, dashed, or dotted line each represents the TTS calculated under the three deterministic MFDs with green time being $\mu_{G}$, $\mu_{G}-\sigma_{G}$, or $\mu_{G}+\sigma_{G}$. Other than the black curves, there are also 1000 scattering points forming the blue curve. Each scatter point is composed of a full disruption recovery process sampled with equal intervals between an initial traffic demand from 1000 to 8000 vehicles. At each time step, a stochastic signal green time is chosen following the normal distribution $G\sim N(14.8,2.5)$ based on real-world data (Genser et al., 2024), leading to an uncertain MFD profile as Fig. 10(b) shows. This is an example of how the congestion is dissipated in the urban road network with an initial disruption of about 7000 vehicles, the green scattering points representing the trip completion at each time step are sampled from every $100^{th}$ point from the total 7200 time steps (seconds) for a clearer and more illustrative plotting.
(a) Demand disruption and TTS
@@
Figure 10: Numerical simulation for demand disruption with stochasticity
Since the blue curve composed from the scattering points closely aligns with the solid curve, it can be inferred that the influence of realistic stochasticity on the MFD is mostly negligible. Nevertheless, an intriguing observation is that, when the disruption demand is relatively low, the blue curve dips slightly below the MFD of the solid curve. However, the blue curve appears to exceed the TTS of the well-defined MFD when the demand is substantial. This may indicate that the recovery process with stochasticity can possibly have a larger second derivative (although two linear curves share the same second derivative of zero but different slopes can yield a similar observation). When we show the distribution of these two curves, as in Fig. 10(c), the TTS with stochasticity has a more concentrated distribution at a lower value while having a marginally longer tail pointing to the right, showing a more left-skewed distribution compared to the one without realistic stochasticity. This can also be validated by calculating the skewness of these two curves. When there is no stochasticity, the skewness is 0.67 while the skewness for the blue curve has a value of 0.70. As a greater skewness indicates a more fragile system, it means by introducing realistic stochasticity, the urban road network becomes even more fragile. It makes particular sense that as per definition, a fragile system should exhibit a much more degraded performance with larger disruptions brought by stochasticity, resulting in poor adaptability to uncertainties.
Likewise, we showcase that supply disruptions can signify the fragile nature of transportation systems as well. With the same simulation environment, instead of the linearly increasing initial disruption demand, now a linearly growing supply disruption magnitude coefficient $r$ from 0 to 0.5 is considered. The simulation of the recovery process from the supply disruption along with 1000 uniformed sampled points between 0 and 0.5 is shown in Fig. 11(a). First of all, as the blue curve with stochasticity lies below the curve from the deterministic MFD with the mean green time, it means that the network with stochasticity has a better performance. This performance improvement can be attributed to the fact that, prior to reaching the maximal capacity, the upper MFD in the dotted curve generated from MoC keeps a larger space from the solid curve MFD profile compared to the distance between the lower MFD in the dashed curve. Therefore, although the likelihood of sampling a trip completion above or below the mean MFD profile is the same, there is a higher probability that the gained value of trip completion will surpass the loss caused by stochasticity. Despite this gain in system performance, when we calculate the skewness of distribution, we get a value of 0.49 for the deterministic MFD and 0.53 for the MFD when considering uncertainty, demonstrating again that stochasticity escalates the fragile response of the road transportation systems. Hence, the transportation systems have been demonstrated to be fragile with numerical simulation and such fragility has been reinforced with stochasticity in this work.
(a) Supply disruption and the TTS
(b) TTS distribution w/wo stochasticity
Figure 11: Numerical simulation for supply disruption with stochasticity
@@ 6 Conclusion
This research systematically demonstrated the fragile nature of road transportation systems with rigorous mathematical analysis and numerical simulation under realistic stochasticity. The mathematical proof comprehends the study of fragility under 1) microscopic - macroscopic, 2) demand disruption - supply disruption, and 3) onset of disruption - recovery from disruption. With essential assumptions regarding the disruption density in comparison to the critical density as well as the base traffic demand, we have validated the fragility of road transportation systems from various perspectives. Furthermore, through a numerical simulation with realistic data, we concluded that real-world stochasticity has a limited impact on the fragile characteristics of the system but contributes to rendering the system even more fragile. As early findings have also pointed out the possibilities of applying MFDs in other transportation modes, the fragility observed in urban road networks may well be extended to various transportation systems. This study aims to offer insights to researchers, emphasizing the fragile characteristics of transportation systems and encouraging the design of antifragile traffic control strategies in the future.
7 CRediT authorship contribution statement
Linghang Sun: Conceptualization, Investigation, Methodology, Visualization, Writing – original draft. Yifan Zhang: Methodology, Visualization, Writing - review & editing. Cristian Axenie: Project administration, Resources, Writing - review & editing. Margherita Grossi: Project administration, Resources, Writing - review & editing. Anastasios Kouvelas: Supervision, Writing - review & editing. Michail A. Makridis: Conceptualization, Methodology, Supervision, Writing - review & editing.
8 Declaration of Competing Interest
This research was kindly funded by the Huawei Munich Research Center under the framework of the Antigones project, with one of our co-authors being employed at the said company. Otherwise, the authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.